Grade 12th PassDifferential Calculus

Find the Range of x4/ (x8+1) .

ATUL KUMAR
13 Years agoGrade 12th Pass

7 Answers

Aman  Bansal
13 Years ago

Dear Atul,

The range is all real except the values of x^8+1 = 0 .

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert


 

ATUL KUMAR
13 Years ago

Dear Aman Bansal

I have already tried finding g(0) then g(1) using g(2) and thus substituting these values but it didn''t helped.

Can you explain how will it work ?

ATUL KUMAR
13 Years ago

Dear Aman Bansal

I have already tried finding g(0) then g(1) using g(2) and thus substituting these values but it didn''t helped.

Can you explain how will it work ?

 

I am asking for th RANGE not DOMAIN .

 

Regards

AK

MUKESH chinna YADAV
13 Years ago

?????????????????

Q.

Find the Range of x4/ (x8+1) .

Sol:-

y=x4/x8+1

y=x4-8+1

y=x-4+1

y=1/x4+1

y=(1+x4)/x4

x4y=1+x4

x4+x4y+1=0

use b2-4ac formula

a=x4,b=x4,c=1

(x4)2-4x4.10<0(<means lessthan or equal to because in my laptop doesn''t support to type less than or equal to )

x8-4x4<0

x8<4x4

x<4x^4/8

x<4x^1/2

x/x^-1/2<0

x^1/2<0

range is (-,0]

ATUL KUMAR
13 Years ago

I don''t understand What should I say for the NONSENSE you have done above. Are you aware of what have you done and to what result you have reached ??

 

y=x4/x8+1

y=x4-8+1 

 

Whats this ?? 

 

x^1/2<0

range is (-∞,0] 

 

HORRIBLE MURDER OF MATHEMATICS ! 

 

Can √x can ever be less than 0 ?? 

 

One more thing range is the Group of possible values of f(x) not x !! 

MUKESH chinna YADAV
13 Years ago

?????the range is all real numbers

Vidyasagar Mohanraj
13 Years ago

For functions like these its best to graph them using calculator to find the range it shows the range is all real numbers